White light containing wavelengths from 4.10E2 nm to7.50E2 nm falls on a grating with 7.8E3 lines/cm. How wide is thefirst-order spectrum on a screen 2.70 m away?

Enter scientific notation as 1.23E4.
Answer #1
d = 1/7.8E3cm = 1.28E-4 cm = 1.28E-6m , L =2.70 m,λ1 = 4.10E2 nm, λ2 =7.50E2 nm
first-orderbright line is at dsinθ =λ
sinθ1 = λ1/d= 0.320, y1 = Ltanθ1 = 0.91 m
sinθ2 = λ2/d = 0.586, y2= Ltanθ2 = 1.95 m
the reqiured width = 1.95 - 0.91 = 1.04 m